# Learning Minimal NAP Specifications for Neural Network Verification

### v2

@@ 2.4. NAP Specifications for Robustness Verification
This subsection illustrates how NAP specifications can be utilized for robustness verification. We discuss how the abstraction states of neurons can serve as signatures of distinct classes with the introduction of the class NAP definition.
@@
In a classification task with the class set $C$, for any class $c\in C$, a class NAP $P^{c}$ is a NAP comprising abstract states outputted by an abstraction function given $X_{c}$, where $X_{c}$ denotes the set of inputs belonging to class $c$. Formally, $P^{c}:=\langle\mathcal{A}(N_{i,l},X_{c})\mid N_{i,l}\in N,\mathcal{A}\in\{\widetilde{\mathcal{A}},\dot{\mathcal{A}}\}\rangle$. The power set of $P^{c}$ is denoted as $\mathcal{P}^{c}$.
Recall that robustness verification can be thought of as proving that no adversarial examples exist in the local neighbourhoods of some reference point $x$. Intuitively, for NAP specifications, it is equivalent to show that no adversarial examples exist in the class NAP $P^{c}$; in other words, inputs exhibiting the NAP $P^{c}$ must be predicted as class $c$. Geng et al. (2023a) argues that class NAPs must satisfy several essential requirements to qualify as NAP specifications. And finding qualified NAP specifications implies the underlying robustness problem is verified. We formally frame these requirements into the following properties:
@@
From a geological perspective, there must be no overlaps between class NAPs. In other words, it is also equivalent to verifying:
@@
We empirically observe that successful models, particularly neural networks that have achieved high accuracy in classification tasks, naturally exhibit this property. Because successful models tend to avoid confusion in predictions, their class NAPs are usually very distinct from each other.
@@
To serve as NAP specifications, class NAPs $P^{c}$ ensure that if an input exhibits it, i.e., $\mathcal{A}(N,x)\preccurlyeq P,\mathcal{A}\in\{\ddot{\mathcal{A}},\dot{\mathcal{A}}\}$, the input must be predicted as the corresponding class $c$. Formally, we have:
(10) $\forall x\in R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad\mathbf{F}_{c}(x)-\mathbf{F}_{k}(x)>0$
@@
(11) $R_{P^{c}}=\{x\mid{\mathcal{A}(N,x)\preccurlyeq P^{c}}\}$
In contrast to canonical $L_{\infty}$ norm balls, class NAPs are more flexible in terms of size and shape. Additionally, there is no need to specify a reference point, since the locations of potential reference points are also encoded by class NAPs. However, it is possible that no class NAP $P^{c}$ in $\mathcal{P}^{c}$ satisfies this property. This can be mitigated by meeting the subsequent weaker property.
@@
Instead of relying solely on class NAPs as specifications, local neighbours can still be employed in conjunction for verification. This hybrid form of specification has several advantages: 1) It narrows down the scope of verifiable regions when no class NAPs can meet the NAP robustness property; 2) NAP constraints essentially fix ReLU states, refining the search space for verification tools; 3) It focuses on the verification of valid test inputs rather than adversarial examples. Formally, this property can be stated as:
(12) $\forall x^{\prime}\in B(x,\epsilon)\bigcap R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad\mathbf{F}_{c}(x)-\mathbf{F}_{k}(x)>0$
@@
(13) $B(x,\epsilon)=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}\quad R_{P^{c}}=\{x^{\prime}\mid{\mathcal{A}(N,x^{\prime})\preccurlyeq P^{c}}\}$
To summarize, we state that a class NAP can serve as a NAP specification if it satisfies either the NAP robustness property or the NAP-augmented robustness property. Clearly, the former property is stronger, and it is possible that we can’t find a class NAP $P^{c}$ in $\mathcal{P}^{c}$ to satisfy this property. Fortunately, we can always find NAPs that satisfy the latter property by narrowing the verifiable region using additional $L_{\infty}$ norm ball specifications. Moreover, NAP specifications can be easily encoded with most of the state-of-the-art verification tools such as Marabou. In this paper, we will not dive into NAP encodings. Interested readers can refer to Geng et al. (2023a) for more details.
@@ 3. The minimal NAP specification problem
In this section, we formulate the problem of learning minimal NAP specifications and present two naive approaches for solving this problem. Since our approaches require interactions with verification tools, we begin by introducing relevant notations describing the relationships between NAPs and verification tools.
Figure 4. A DAG representing all possible NAPs, with green nodes signifying they can serve as a specification, while red nodes indicate they cannot. Within each node, a black circle means the corresponding neuron is abstracted by $\widetilde{\mathcal{A}}$, whereas the white circle means it is abstracted by $\dot{\mathcal{A}}$. Thus, the top node represents the coarsest NAP, while the bottom node entails the most refined one. Our goal is to find the NAP that resides in the green nodes while its size is minimal.
@@ 3.1. Problem Formulation
Let $(\mathcal{P}^{c},\preccurlyeq)$ be a partially ordered set corresponding to a family of class NAPs regarding some class $c\in C$. For simplicity, we omit the superscript $c$ and refer to class NAPs simply as NAPs when the context is clear. We assume we have access to a verification tool, $\mathcal{V}:\mathcal{P}\to\{0,1\}$, which maps a class NAP $P\in\mathcal{P}$ to a binary set. Here, $\mathcal{V}(P)=1$ denotes a successful verification of the underlying robustness query, while 0 indicates the presence of an adversarial example. From an alternative perspective, $\mathcal{V}(P)=1$ also signifies that $P$ is a NAP specification, i.e., it satisfies NAP(-augmented) robustness properties; whereas $\mathcal{V}(P)=0$ implies the opposite.
It is not hard to see that $\mathcal{V}$ is monotone with respect to the NAP family $(\mathcal{P},\preccurlyeq)$. Given $P\preccurlyeq P^{\prime}$ and $\mathcal{V}(P^{\prime})=1$, it follows that $\mathcal{V}(P)=1$. However, given $P\preccurlyeq P^{\prime}$ and $\mathcal{V}(P)=1$, we cannot determine $\mathcal{V}(P^{\prime})$. In other words, refining a NAP (by increasing the number of neurons abstracted to $\mathbf{0}$ or $\mathbf{1}$) can only enhance the likelihood of successful verification of the underlying robustness query.
@@
Given a family of NAPs $\mathcal{P}$ and a verification tool $\mathcal{V}$, the minimal NAP specification problem is to find a NAP $P$ such that
$\underset{P\in\mathcal{P},\mathcal{V}(P)=1}{\arg\min}|P|$
where $|P|$ is defined as the size of $P$, representing the number of neurons that are abstracted to the $\mathbf{0}$ or $\mathbf{1}$, i.e., |$|\{N_{i,l}\mid P_{i,l}=\mathbf{0}\text{ or }\mathbf{1}\}|$. The (largest) minimal NAP specification’s size is denoted as $s$.
@@
When $P$ is minimal, it implies that for any NAP $P^{\prime}$ that is strictly more coarse than $P$, $\mathcal{V}(P^{\prime})=0$. Formally, this can be expressed as: $\forall P\preccurlyeq P^{\prime},P^{\prime}\neq P:\mathcal{V}(P^{\prime})=0$. Thus, there could exist multiple minimal NAP specifications. In such cases, we only need to choose one of them. On the other hand, it is possible that even the most refined NAPs cannot verify the robustness query. In such cases, we claim that no minimal NAP specifications exist. Additionally, since the computation using verification tools is usually expensive, we are interested in methods that efficiently find a minimal NAP specification, i.e., minimizing the number of calls to $\mathcal{V}$.
We present two naive approaches to solving the problem. Since these approaches involve refining and coarsening the NAP, we will first formally define these two actions.
Refine and Coarsen Action of NAP
Recall our definition of NAP, the coarsest NAP is the one using $\dot{\mathcal{A}}$ to abstract each neuron in $N$. We denote this NAP as $\dot{P}:=\langle\dot{\mathcal{A}}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N\rangle$. $\dot{P}$ is the smallest NAP with $|\dot{P}|=0$. In addition, we define the most refined NAP as the one that applies $\widetilde{\mathcal{A}}$ to abstract each neuron in $N$. We denote it as $\widetilde{P}:=\langle\widetilde{\mathcal{A}}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N\rangle$. $\widetilde{P}$ is the largest NAP with size with $|\widetilde{P}|\leq|N|$. Clearly, $\widetilde{P}\preccurlyeq\dot{P}$. Given any NAP $P$, if we want to refine $P$ through a specific neuron $N_{i,l}$, we apply the $\widetilde{\mathcal{A}}$ function to $N_{i,l}$. We denote this refinement action as $\widetilde{\Delta}(N_{i,l})$. This action will either increase or leave unchanged $|P|$. Similarly, we denote the coarsen action as $\dot{\Delta}(N_{i,l})$. This action will either decrease or leave unchanged $|P|$. We present the semantics of these actions as follows:
[…]
$P\vdash\cdot\Downarrow P$
[…]
where $P,P^{\prime},P^{\prime\prime}$ are from the NAP family $\mathcal{P}$, and $N$ is the underlying neural network. $\Delta_{1}$ and $\Delta_{2}$ are any two sequences of refine and coarsen actions.
3.2. The Refine Approach
Conceptually, the Refine approach iteratively increases the number of refined neurons in NAP $P$ until $\mathcal{V}(P)=1$, i.e., $P$ is able to prove the underlying robustness query. In other words, we gradually increase the size parameter $k$ and iterate over each NAP $P$ of size $k$ to check if $\mathcal{V}(P)=1$, as illustrated in Algorithm 1. To determine if a solution to the problem exists, we first check if the most refined NAP can succeed in verification. We proceed to iterative refinement only if $\mathcal{V}(\widetilde{P})=1$. However, the algorithm is not efficient and requires $12^{|N|}-$ 1 calls to $\mathcal{V}$ in the worst case, as proven in Theorem 3.2. Please refer to the proof in Appendix A. Therefore Refine is only practical when the search space of the NAP family $\mathcal{P}$ is small.
Theorem 3.2.
3.3. The Coarsen Approach
@@
In contrast to the Refine approach, the Coarsen approach starts from the most refined NAP and then gradually coarsens it. We first check if the problem is well-defined by verifying if the most refined NAP $\widetilde{P}$ succeeds in verification. Then, for each neuron, we attempt to coarsen it using $\dot{\mathcal{A}}$; if the resulting NAP no longer verifies the query, we refine it back using $\widetilde{\mathcal{A}}$; otherwise, we keep the coarsened NAP. We describe the above procedure in Algorithm 2. The algorithm could require $|N|$ calls to $\mathcal{V}$ in the worst case, as proven in Theorem 3.3. Please refer to the proof in Appendix A.

### v4

@@ 2.4. NAP Specifications for Robustness Verification
The previously introduced NAP definition primarily serves as a conceptual tool. To function as a specification for robustness verification, the regions outlined by NAPs ($R_{P}$) should cover a significant amount of data from a specific class. We demonstrate how to achieve this using the $\widetilde{\mathcal{A}}$ function.
The statistical $\widetilde{\mathcal{A}}$ function
The binary $\ddot{\mathcal{A}}$ function is limited to processing a single input and encounters challenges when multiple inputs are involved, as two distinct inputs may lead to conflicting abstraction states for a neuron; for instance, a neuron may be deactivated for input $x_{1}$ but activated for another input $x_{2}$. This can be problematic when attempting to learn a general NAP for an entire class of inputs. To address this issue, we approach it statistically by introducing the $\widetilde{\mathcal{A}}$ function, defined as follows:
where $\delta$ is a real number from $[0,1]$, and $X$ represents a set of inputs, i.e., $X:=\{x_{1},x_{2},\ldots,x_{n}\}$. Since datasets often contain noisy data or challenging instances that the model cannot predict accurately, we introduce the parameter $\delta$ to accommodate standard classification settings in which Type I and Type II errors are non-negligible. Intuitively, the introduction of $\delta$ allows multiple inputs $x_{1},...,x_{n}$ to vote on a neuron’s state. For instance, when $\delta$ is set to 0.99, then 99% or more of the inputs must agree that a neuron is activated for the neuron to be in the $\mathbf{1}$ state. It is worth mentioning that the statistical $\widetilde{\mathcal{A}}$ function is equivalent to the method described in Geng et al. (2023).
@@
In a classification task with the class set $C$, for any class $c\in C$, a class NAP $P^{c}$ is a NAP comprising abstract states outputted by an abstraction function given $X_{c}$, where $X_{c}$ denotes the set of inputs belonging to class $c$. Formally, $P^{c}:=\langle\mathcal{A}(N_{i,l},X_{c})\mid N_{i,l}\in N\rangle$, where $\mathcal{A}$ is either $\widetilde{\mathcal{A}}$ or $\dot{\mathcal{A}}$, chosen for each neuron. The power set of $P^{c}$ is denoted as $\mathcal{P}^{c}$.
Robustness verification involves proving that no adversarial examples exist in the local neighborhood of a reference point $x$. For NAP specifications, this translates to showing that no adversarial examples exist in class NAP $P^{c}$; in other words, inputs exhibiting $P^{c}$ must be classified as $c$. Geng et al. (2023) argues that class NAPs must meet several key requirements to qualify as NAP specifications, and finding such NAPs effectively verifies the underlying robustness problem. We formalize these requirements into the following properties:
@@
From a geometric perspective, there must be no overlaps between class NAPs. In other words, it is also equivalent to verifying:
@@
To serve as NAP specifications, class NAPs $P^{c}$ ensure that if an input exhibits it, i.e., $P\preccurlyeq\ddot{\mathcal{A}}(N,x)$, the input must be predicted as the corresponding class $c$. Formally, we have:
(10) $\forall x\in R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad F_{c}(x)-F_{k}(x)>0$
@@
(11) $R_{P^{c}}=\{x\mid{P^{c}\preccurlyeq\ddot{\mathcal{A}}(N,x)}\}$
@@
Rather than relying exclusively on class NAPs as specifications, local neighbors can also be used in combination for verification. This hybrid approach offers several advantages: 1) It narrows the scope of verifiable regions when class NAPs alone cannot satisfy the NAP robustness property; 2) NAP constraints essentially lock ReLU states, thereby refining the search space for verification tools; 3) It emphasizes verifying valid test inputs, while still covering broader and more flexible verifiable regions than those defined by $L_{\infty}$ norm ball specifications alone. Formally, this property can be expressed as:
(12) $\forall x^{\prime}\in B(x,\epsilon)\bigcap R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad F_{c}(x)-F_{k}(x)>0$
@@
(13) $B(x,\epsilon)=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}\quad R_{P^{c}}=\{x^{\prime}\mid{P^{c}\preccurlyeq\ddot{\mathcal{A}}(N,x^{\prime})}\}$
To summarize, we state that a class NAP can serve as a NAP specification if it satisfies either the NAP robustness property or the NAP-augmented robustness property. Clearly, the former property is stronger, and it is possible that we can’t find a class NAP $P^{c}$ in $\mathcal{P}^{c}$ to satisfy this property. Fortunately, we can always find NAPs that satisfy the latter property by narrowing the verifiable region with additional $L_{\infty}$ norm ball specifications.
@@ 3. The minimal NAP specification problem
In this section, we formulate the problem of learning minimal NAP specifications and present both a deterministic and statistical approach to address it. As our methods involve interactions with verification tools, we first introduce the relevant notations.
@@ 3.1. Problem Formulation
Let $(\mathcal{P}^{c},\preccurlyeq)$ be a partially ordered set corresponding to a family of class NAPs regarding some class $c\in C$. For simplicity, we omit the superscript $c$ and refer to class NAPs simply as NAPs when the context is clear. We assume access to a verification tool, $\mathcal{V}:\mathcal{P}\to\{0,1\}$, which maps a class NAP $P\in\mathcal{P}$ to a binary set. Here, $\mathcal{V}(P)=1$ denotes a successful verification of the underlying robustness query, while 0 indicates the presence of an adversarial example. From an alternative perspective, $\mathcal{V}(P)=1$ also signifies that $P$ is a NAP specification, i.e., it satisfies NAP(-augmented) robustness properties; whereas $\mathcal{V}(P)=0$ implies the opposite.
It is not hard to see that $\mathcal{V}$ is monotone with respect to the NAP family $(\mathcal{P},\preccurlyeq)$. Given $P^{\prime}\preccurlyeq P$ and $\mathcal{V}(P^{\prime})=1$, it follows that $\mathcal{V}(P)=1$. However, given $P^{\prime}\preccurlyeq P$ and $\mathcal{V}(P)=1$, we cannot determine $\mathcal{V}(P^{\prime})$. In other words, refining a NAP (by increasing the number of neurons abstracted to $\mathbf{0}$ or $\mathbf{1}$) can only enhance the likelihood of successful verification of the underlying robustness query.
@@
Given a family of NAPs $\mathcal{P}$ and a verification tool $\mathcal{V}$, the minimal NAP specification problem is to find a NAP $P$ such that:
$\forall P^{\prime}\preccurlyeq P,\,P^{\prime}\neq P\implies\mathcal{V}(P^{\prime})=0$
That is, when $P$ is minimal, any NAP $P^{\prime}$ that is strictly coarser than $P$ will result in $\mathcal{V}(P^{\prime})=0$. The size of $P$, denoted by $|P|$ or $s$, represents the number of neurons abstracted to $\mathbf{0}$ or $\mathbf{1}$, i.e., |$|\{N_{i,l}\mid P_{i,l}=\mathbf{0}\text{ or }\mathbf{1}\}|$.
@@
It is important to note that ”minimal” refers to the level of abstraction, not the size of the NAP. Since $(\mathcal{P}^{c},\preccurlyeq)$ is a partially ordered set, multiple minimal NAP specifications may exist, rather than a single global minimum. In such cases, selecting any one of the minimal NAPs is sufficient. However, it is possible that even the most refined NAPs cannot verify the robustness query, in which case no minimal NAP specification exists. Given the computational cost of using verification tools, we are particularly interested in methods that efficiently find a minimal NAP specification, minimizing the number of calls to $\mathcal{V}$.
3.2. Conservative Bottom Up Approach
We introduce Coarsen, a method that starts with the most refined NAP and gradually coarsens it to learn minimal NAP specifications. Before proceeding, we define the relevant notations.
Given a class input $X_{c}$, the coarsest NAP is defined as the one that uses $\dot{\mathcal{A}}$ to abstract each neuron in $N$, denoted as $\dot{P}:=\langle\dot{\mathcal{A}}(N_{i,l},X_{c})\mid N_{i,l}\in N\rangle$. This NAP is the smallest possible in size, with $|\dot{P}|=0$. Conversely, the most refined NAP applies $\widetilde{\mathcal{A}}$ to abstract each neuron, denoted as $\widetilde{P}:=\langle\widetilde{\mathcal{A}}(N_{i,l},X_{c})\mid N_{i,l}\in N\rangle$ or $\widetilde{\mathcal{A}}(N)$. This NAP is the largest in size, with $|\widetilde{P}|\leq|N|$. Clearly, we have $\dot{P}\preccurlyeq\widetilde{P}$.
To refine any NAP $P$ through a specific neuron $N_{i,l}$, we apply the $\widetilde{\mathcal{A}}$ function, denoting this refinement as $\widetilde{\Delta}(N_{i,l})$. This action may either increase or leave $|P|$ unchanged. Conversely, we denote the coarsen action as $\dot{\Delta}(N_{i,l})$, which will either decrease or leave $|P|$ unchanged.
@@ 3.2. Conservative Bottom Up Approach
The Coarsen approach starts by verifying if the most refined NAP $\widetilde{P}$ successfully passes verification. For each neuron, we attempt to coarsen it using $\dot{\mathcal{A}}$. If the resulting NAP no longer verifies the query, we revert to refining it back using $\widetilde{\mathcal{A}}$; otherwise, we retain the coarsened NAP. The procedure is detailed in Algorithm 1, which may require up to $|N|$ calls to $\mathcal{V}$ in the worst case, as shown in Theorem 3.2. Please refer to the proof in Appendix D.
Input: The neural network $N$
Output: A minimal NAP specification $P$
1 Function Coarsen(N)
2       $P\leftarrow\widetilde{\mathcal{A}}(N)$
3       if $\mathcal{V}(P)==0$ then
             return $None$ ;
              /* Return None if the most refined NAP fails verification */
4
5       else
6             for $N_{i,l}$ in $N$ do
                   $P\leftarrow\dot{\Delta}(N_{i,l})$ ;
                    /* Try to coarsen(abstract) this neuron using $\dot{\Delta}$/
7                   if $\mathcal{V}(P)==0$ then
                         $P\leftarrow\widetilde{\Delta}(N_{i,l})$ ;
                          /* Refine the neuron back if the verification fails */
8
9
10            return $P$;
11
Algorithm 1 Coarsen